OpenStudy (anonymous):

integral dx/8+2x^2 from 0 to 2. I only need help changing the limits of integration

OpenStudy (anonymous):

when you say x=0... where do you plug that 0 into? I defined x=3tantheta

OpenStudy (anonymous):

take 8 common from denom....1/8 dx/1+(x/2)^2 then use integration of tanx

OpenStudy (anonymous):

\[\int\limits_{0}^{2}\frac{ dx }{ 8+2x ^{2} }=\frac{ 1 }{ 2 } \int\limits_{0}^{2} \frac{ dx }{ 4+x ^{2} }, put x=2 \tan \theta \]

OpenStudy (anonymous):

for solving yes.. i just plugged into the basic integral formula.. I am talking about doing it by trig sub and changing the limits. I could put a problem up that cannot be substituted .. this was just the one I was working on and I realize I do not understand how to change the limits of integration if we are going from theta back to x

OpenStudy (anonymous):

i mean if we are not going back to x

OpenStudy (anonymous):

in the sol manual the changed the limits from 0 to pi/4 and I am wondering how i get there.

OpenStudy (anonymous):

\[when x=0,\theta=0,when x=2,\tan \theta=1,\theta=\frac{ \pi }{ 4 }\]

OpenStudy (anonymous):

you plug it back in for theta? thus theta = arctan x/2 .. subbing original limits?

OpenStudy (anonymous):

theta=arctan0/2 = 0 ... and 2/2 is pi/4... woo I think I understand!!!

OpenStudy (anonymous):

thank you for your help.

OpenStudy (anonymous):

just integrate . \[=\frac{ 1 }{ 2 } \int\limits_{0}^{\frac{ \pi }{ 4 }}\frac{ 2\sec ^{2}\theta d \theta }{4\left( 1+\tan ^{2}\theta \right) }\]